The equilibrium constant for the reaction

H₂(g) + S(s) ⇔H₂S(g) is 20.0 at 925K and 10.25 at 1000K, respectively. To calculate the standard enthalpy of the reaction.


1
Expert's answer
2021-01-15T06:46:25-0500

This can be done by using formula from Vant Hoff eqation, from which we an write

log(k2\k1) = ("\Delta"H\2.303xR)x(1\T1 - 1\T2)where k2 = 10.25 , k1 = 20.0 , T1 = 925 K , T2 = 1000k , R = 8.314 J mol-1K-1

"\Delta"H = standard enthalpy chnage

or, log(10.25\20.0)= log(0.5125) = -0.29 = ("\Delta"H\2.303xR)x(1\T1 - 1\T2)

-0.29x2.303x8.314 Jmol-1k-1 = "\Delta"H x (1\925 - 1\1000) K-1

or, 5.55 Jmol-1 = "\Delta"H x (0.00108 - 0.001) = "\Delta"H x 8 x 10-5

or , "\Delta H = (5.55"\8)x105 Jmol-1 = 69375 Jmol-1 = 69.375 kJ mol-1 (ans)